### v1

@@ 3.2 Gaussian Process Regression
The surrogate model proposed in this study (see subsection 3.4) employs GPR to perform regression on variables in the latent space derived from the $\beta$-VAE. Using flight conditions as input data points, $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n}\}\in\mathcal{P}$ (the parameter space), where each $\mathbf{p}_{i}=[M_{i},\alpha_{i}]$, the GPR model predicts the set of $d$ latent variables within the latent space, $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{n}\}$, where each $\mathbf{y}_{i}=[\mu_{i1},\ldots,\mu_{id}]$ represents the latent variables associated with $\mathbf{p}_{i}$. The choice of GPR is justified by its effectiveness in performing non-linear regressions, offering high accuracy and flexibility, thereby ensuring robust and reliable models. An overview of GPR theory is presented below. For a detailed mathematical analysis, readers are referred to comprehensive literature on the topic [53, 54, 55].
A Gaussian process (GP) is a collection of random variables, any finite number with a joint Gaussian distribution. It is characterized by its mean function, $m(\mathbf{p})=\mathbb{E}[f(\mathbf{p})]$, and covariance function, $k(\mathbf{p},\mathbf{p}^{\prime})=\mathbb{E}[(f(\mathbf{p})-m(\mathbf{p}))(f(\mathbf{p}^{\prime})-m(\mathbf{p}^{\prime}))]$. Thus, the classic definition of a GP is given by:
@@ 3.4 Surrogate modelling
The kernel selection for GPR is based on the geometry of the latent space and its relationship with the parameter space. A linear kernel is proposed, supported by the observed linear solid relationship between the parameters and latent variables. This linear kernel effectively models direct relationships, providing a solid basis for regression tasks. Additionally, the linear kernel is combined with the Matérn 3/2 kernel to capture the data set’s non-linearities. The Matérn 3/2 kernel is well-suited for modeling smooth functions with significant non-linear variations, making it appropriate for capturing the inherent complexities and variabilities in the data. Hence, the kernel is given by:
@@
Combining these kernels multiplicatively, the GPR model captures both linear and non-linear dependencies, enhancing its predictive capability and ensuring a more accurate and robust representation of the relationships between flight conditions and latent space variables.
@@ 4 Latent space analysis
To further assess autoencoder performance, the physical interpretability of the latent space is crucial. The latent variables are ordered based on their contribution to the global reconstruction loss of the autoencoder, allowing the identification of the most relevant ones. Hence, the physical features contained in each LV can be assessed. Figure 4 shows the distribution of data in the latent space for both $\beta$-VAE (a) and PCA+$\beta$-VAE (b) models, focusing on the plane formed by the two most relevant latent variables. The color gradient represents the Mach number, while the size of each data point indicates $\alpha$. This visualization provides insights into how the latent variables capture the underlying flight conditions. For both $\beta$-VAE and PCA+$\beta$-VAE models, there is a perceptible distribution of data points that correlates with the flight conditions $[M,\alpha]$, which are well-represented in the latent space, indicating that the models effectively capture these critical aerodynamic parameters. For instance, in the case of $\beta=0.008$ for the $\beta$-VAE model, there is a clear correlation between the flight condition parameters and the latent variables $\mu_{1}$ and $\mu_{2}$: Mach number increasing from bottom to top and $\alpha$ from left to right. This result is noteworthy because the fully unsupervised autoencoder (no labels of the flight conditions are provided) learns to differentiate among flight conditions based on their associated pressure fields, leading to a smooth latent space ideal for the subsequent regression task. This behavior is consistent across different latent space dimensions and $\beta$ values. The PCA+$\beta$-VAE model shows a slightly more structured and stable distribution across different $\beta$ values, suggesting that the PCA pre-processing stage enables a better converged $\beta$-VAE training.
@@
The reconstruction error in Figure 3 is directly linked to the data distribution in the latent space (Figure 4), strongly influenced by $\beta$. For minimal $\beta$ values (close to 0), both models show a well-organized latent space with distinct clustering according to flight conditions. As $\beta$ increases, the clustering becomes less distinct, particularly at $\beta=1$, where the regularization effect of the KL divergence term dominates, leading to a less informative latent space. The non-linear relationship between $M$ and $\alpha$ is well captured at low $\beta$, but as $\beta$ increases and latent space orthogonality is promoted, the relationship between the parameter space $[M,\alpha]$ and latent variables $[\mu_{1},\mu_{2}]$ becomes unclear. On the contrary, as the latent space dimension increases (from 2 to 6), the data distribution in the latent space remains relatively consistent, particularly for smaller values of $\beta$, which suggests that the most relevant features are adequately captured within the first two latent variables, supporting the decision to focus on this lower-dimensional representation. The remaining dimensions are energetically poor, with no significant information content. This fact is especially evident when PCA pre-processing is introduced, as most latent spaces become visually sorted compared to the $\beta$-VAE model, illustrating the dimensionality reduction effect discussed in Figure 3.
@@ 5.1 Regression within the latent space
The regressor tends to fail more frequently at points near or on the boundaries of the parameter space, especially at high angles of attack (the right side of the latent space). This pattern suggests that the regressor struggles more in these regions, possibly due to the higher complexity and non-linearity of the aerodynamic responses, such as flow separation or displacement of shockwaves. Additionally, these boundary data points in the latent space have fewer neighbors, reducing the local information available for the regressor and making accurate predictions more challenging.
@@ 5.3 Fine-tuning process: towards $\beta$-independent surrogate model
Table 3: erformance indicators of the fine-tuned models. They include MAE, RMSE, and R2 scores, which have been computed using the testing dataset.
@@ 6 Conclusions
An analysis of the application of $\beta$-variational autoencoders combined with regression based on Gaussian processes has been conducted to predict pressure fields over a wing subjected to transonic flight conditions. The complexity of this three-dimensional flow allows us to evaluate the capability of data-driven models to predict non-linearities in pressure fields due to shockwaves, high angles of attack, and the effects of high load factor conditions.
@@ Acknowledgments
The dataset is subjected to the intellectual property of AIRBUS and cannot be shared, unless previous approval. The software developed for the surrogate framework will be available on GitHub upon publication.

### v2

@@ 3.2 Gaussian Process Regression
The surrogate model proposed in this study (see subsection 3.4) employs GPR to perform regression on variables in the latent space derived from the $\beta$-VAE. Using flight conditions as input data points, $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n}\}\in\mathcal{P}$ (the parameter space), where each $\mathbf{p}_{i}=[M_{i},\alpha_{i}]$, the GPR model predicts the set of $d$ latent variables within the latent space, $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{n}\}$, where each $\mathbf{y}_{i}=[\mu_{i1},\ldots,\mu_{id}]$ represents the latent variables associated with $\mathbf{p}_{i}$. The choice of GPR is justified by its effectiveness in performing nonlinear regressions, offering high accuracy and flexibility, thereby ensuring robust and reliable models. An overview of GPR theory is presented below. For a detailed mathematical analysis, readers are referred to comprehensive literature on the topic [55, 56, 57].
A Gaussian process (GP) is a collection of random variables, any finite number with a joint Gaussian distribution. It is characterized by its mean function, $m(\mathbf{p})=\mathbb{E}[f(\mathbf{p})]$, and covariance function, $k(\mathbf{p},\mathbf{p}^{\prime})=\mathbb{E}[(f(\mathbf{p})-m(\mathbf{p}))(f(\mathbf{p}^{\prime})-m(\mathbf{p}^{\prime}))]$, where $\mathbf{p}$ is the input vector, and $\mathbb{E}$ is the expected value. Thus, the classic definition of a GP is given by:
@@ 3.4 Surrogate modelling
The kernel selection for GPR is based on the geometry of the latent space and its relationship with the parameter space. A linear kernel is proposed, supported by the observed linear solid relationship between the parameters and latent variables. This linear kernel effectively models direct relationships, providing a solid basis for regression tasks. Additionally, the linear kernel is combined with the Matérn 3/2 kernel to capture the data set’s nonlinearities. The Matérn 3/2 kernel is well-suited for modeling smooth functions with significant nonlinear variations, making it appropriate for capturing the inherent complexities and variabilities in the data. Hence, the kernel is given by:
@@
Combining these kernels multiplicatively, the GPR model captures both linear and nonlinear dependencies, enhancing its predictive capability and ensuring a more accurate and robust representation of the relationships between flight conditions and latent space variables.
@@ 4 Latent space analysis
To further assess autoencoder performance, the physical interpretability of the latent space is crucial. The latent variables are ordered based on their contribution to the global reconstruction loss of the autoencoder, allowing the identification of the most relevant ones. To rank the latent variables, the model is evaluated by setting one mode to zero at a time. Modes that result in the highest error when set to zero are considered to contain the most information and are subsequently ordered [44, 45]. Hence, the physical features contained in each LV can be assessed. Figure 4 shows the distribution of data in the latent space for both $\beta$-VAE (a) and PCA+$\beta$-VAE (b) models, focusing on the plane formed by the two most relevant latent variables. The color gradient represents the Mach number, while the size of each data point indicates $\alpha$. This visualization provides insights into how the latent variables capture the underlying flight conditions. For both $\beta$-VAE and PCA+$\beta$-VAE models, there is a perceptible distribution of data points that correlates with the flight conditions $[M,\alpha]$, which are well-represented in the latent space, indicating that the models effectively capture these critical aerodynamic parameters. For instance, in the case of $\beta=0.008$ for the $\beta$-VAE model, there is a clear correlation between the flight condition parameters and the latent variables $\mu_{1}$ and $\mu_{2}$: Mach number increasing from bottom to top and $\alpha$ from left to right. This result is noteworthy because the fully unsupervised autoencoder (no labels of the flight conditions are provided) learns to differentiate among flight conditions based on their associated pressure fields, leading to a smooth latent space ideal for the subsequent regression task. This behavior is consistent across different latent space dimensions and $\beta$ values. The PCA+$\beta$-VAE model shows a slightly more structured and stable distribution across different $\beta$ values, suggesting that the PCA pre-processing stage enables a better converged $\beta$-VAE training.
@@
The reconstruction error in Figure 3 is directly linked to the data distribution in the latent space (Figure 4), strongly influenced by $\beta$. For minimal $\beta$ values (close to 0), both models show a well-organized latent space with distinct clustering according to flight conditions. As $\beta$ increases, the clustering becomes less distinct, particularly at $\beta=1$, where the regularization effect of the KL divergence term dominates, leading to a less informative latent space. The nonlinear relationship between $M$ and $\alpha$ is well captured at low $\beta$, but as $\beta$ increases and latent space orthogonality is promoted, the relationship between the parameter space $[M,\alpha]$ and latent variables $[\mu_{1},\mu_{2}]$ becomes unclear. On the contrary, as the latent space dimension increases (from 2 to 6), the data distribution in the latent space remains relatively consistent, particularly for smaller values of $\beta$, which suggests that the most relevant features are adequately captured within the first two latent variables, supporting the decision to focus on this lower-dimensional representation. The remaining dimensions are energetically poor, with no significant information content. This fact is especially evident when PCA pre-processing is introduced, as most latent spaces become visually sorted compared to the $\beta$-VAE model, illustrating the dimensionality reduction effect discussed in Figure 3.
@@ 5.1 Regression within the latent space
The regressor tends to fail more frequently at points near or on the boundaries of the parameter space, especially at high angles of attack (the right side of the latent space). This pattern suggests that the regressor struggles more in these regions, possibly due to the higher complexity and nonlinearity of the aerodynamic responses, such as flow separation or displacement of shockwaves. Additionally, these boundary data points in the latent space have fewer neighbors, reducing the local information available for the regressor and making accurate predictions more challenging.
@@ 5.3 Fine-tuning process: towards $\beta$-independent surrogate model
Table 3: Performance indicators of the fine-tuned models. They include MAE, RMSE, and R2 scores, which have been computed using the testing dataset.
@@ 6 Conclusions
An analysis of the application of $\beta$-variational autoencoders combined with regression based on Gaussian processes has been conducted to predict pressure fields over a wing subjected to transonic flight conditions. The complexity of this three-dimensional flow allows us to evaluate the capability of data-driven models to predict nonlinearities in pressure fields due to shockwaves, high angles of attack, and the effects of high load factor conditions.
@@ Acknowledgments
The dataset is subjected to the intellectual property of AIRBUS and cannot be shared, unless previous approval. The software for the surrogate has been primarily developed using PyTorch and scikit-learn frameworks. The codes used for this work are available on GitHub at https://github.com/FluidMechanicsInta/Towards-aerodynamic-surrogate-modeling-based-on-beta-variational-autoencoders.